OpenStudy (anonymous):

The circumference of a coin is 12TT cm. What is the radius? What is the diameter?

OpenStudy (campbell_st):

using the formula for circumference \[C = 2 \pi r\] substitute your information \[12\pi = 2\pi r\] and solve for r

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the next step

OpenStudy (campbell_st):

divide both sides of the equation by 2 pi...

OpenStudy (anonymous):

how do I do that

OpenStudy (campbell_st):

well its